OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4. y = -2x + 7 y = 2x - 3 y = 2x - 7 y = -2x + 3

OpenStudy (anonymous):

If the line is parallel, it will have the same slope as the original equation, therefore you can get rid of A and D. To finish solving, you write an equation that looks like this:\[-5=2(1)+b \] Then solve for b. -5=2+b -7=b So the answer is y=2x-7.

OpenStudy (anonymous):

thank you